Q1. Acetylene[C2H2] burns in air forming carbon dioxide and water vapour. Calcu;late the volume of air reqired to completely burn 50cm3 of acetylene [assume air contains 20% of axygen
1
Expert's answer
2013-02-13T11:51:39-0500
Reaction of burning of acetylene:

50cm^3 V
2HC=CH + 3O2 = 4CO2 + 2H2O
2 3

Volume is next:

V= 50 * 3 /2 = 75 cm3 of Oxygen required for this burning.

If 75 cm3 is 20 %
x is 100%

x= 375 cm^3 of air.